Q: 2 A firm faces the production function Q = 12K0.4 L0.4 and can buy the inputs K and L at prices per unit of £40 and £5 respectively. If it has a budget of £800 what combination of K and L should it use in order to produce the maximum possible output? Show graphically as well.



1
Expert's answer
2021-05-17T10:22:15-0400

The problem is to maximize the function "Q = 12K^{0.4}L^{0.4}" subject to the budget constraint

"40K + 5L = 800"

The theory of the firm tells us that a firm is optimally allocating a fixed budget if the last £1 spent on each input adds the same amount to output, i.e. marginal product over price should be equal for all inputs. This optimization condition can be written as

"\frac{MP_K}{P_K} = \frac{MP_L}{P_L}"

The marginal products can be determined by partial differentiation:

"MP_K = \frac{dQ}{dK} = 4.8K^{-0.6}L^{0.4} \\

MP_L = \frac{dQ}{dL} = 4.8K^{0.4}L^{-0.6} \\

4.8K^{-0.6}L^{0.4} = 4.8K^{0.4}L^{-0.6}"

Dividing both sides by 4.8 and multiplying by 40 gives:

"K^{-0.6}L^{0.4}=8K^{0.4}L^{-0.6}"

Multiplying both sides by "K^{0.6}L^{0.6}" gives:

"L = 8K \\

40K + 5 \times 8K = 800 \\

40K + 40K = 800 \\

80K=800 \\

K = 10 \\

L = 8 \times 10 = 80"